# Unveiling a 36 Billion Solar Mass Black Hole at the Centre of the Cosmic Horseshoe Gravitational Lens

### v1

@@ 3.2 Dynamical modelling
and applying the boundary condition $\nu\overline{v^{2}_{r}}=0$ as $r\rightarrow 0$, the Jeans equations 7 simplify to:
@@
For this purpose, we employed the Jeans Axisymmetric Multi-Gaussian Expansion (JAM) method (Cappellari, 2008, 2020), as implemented in the Jampy software, to perform the stellar dynamical modelling. This approach assumes that the galaxy’s mass distribution can be parameterised as a sum of concentric elliptical Gaussians (see Section 3.3).
@@ 3.3 Multi-Gaussian Expansion (MGE)
The stellar light and mass component is modelled using the MGE method (Emsellem et al., 1994; Cappellari, 2002), which parametrises the stellar surface brightness as a sum of concentrict elliptical Gaussians with the same orientation. If $I(x^{\prime},y^{\prime})$ represents the stellar surface brightness, its MGE parametrisation is given by
@@
Here, $M_{j}=\Upsilon_{\star}L_{j}$ represents the mass of the $j^{\text{th}}$ Gaussian component, with $L_{j}$ luminosity and $\sigma_{j}$ dispersion, as defined in Equation 12. The deprojected axial ratio, $q_{j}$, is given by
@@ 3.4 Mass profile
We described the mass profile of the main deflector using the MGE method, assuming a multicomponent mass model composed by (i) a stellar mass component; (ii) a DM halo component; and (iii) an additional central mass concentration representing a SMBH.
@@
where $\rho_{s}$ is a characteristic density at the scale radius $r_{s}$, and $\gamma_{\text{DM}}$ is the inner density slope that allows the profile to be cuspier ($\gamma_{\text{DM}}>1$) or cored ($\gamma_{\text{DM}}=0$). For $\gamma_{\text{DM}}=1$, the profile reduces to the classical NFW (Navarro et al., 1997). To include this halo contribution in the dynamical model, we followed Cappellari et al. (2013) and also parametrised the DM component using the MGE method. Finally, the SMBH is modelled as an additional Gaussian with a small scale radius ($\sigma=0.01\arcsec$).
@@ 3.5 Joint modelling
We break the modelling process into several stages, so as to make it more tractable
@@
    Dynamical Model Fitting: We begin by fitting only the dynamical model, but with the addition of the prior term for  $M_{\text{Ein}}$ and the conjugate regions of s1. This provides an initial estimate of the mass parameters.
@@
    Mass Parameter Resampling: After obtaining the source grid, we resample the mass parameters using the full likelihood, with s1 reconstructed based on the highest-likelihood model from the previous chain. This step improves the estimate of the mass parameters.
@@
    Source Plane Resampling: Using the highest-likelihood model, we resample the parameters of the source plane. The source is reconstructed in the source plane using a brightness-adaptive Voronoi mesh grid, with Natural Neighbour interpolation (Sibson, 1981) and a regularisation scheme that adapts the degree of smoothing based on the source’s surface brightness. This ensures that brighter regions are reconstructed at higher resolution, while regions with lower SNR are more regularised. In PyAutoLens notation, this corresponds to a KMeans mesh grid, with VoronoiNN pixelisation, and AdaptiveBrightnessSplit regularisation.
@@
    Final Resampling of the Lens Macro Model: Finally, using the source grid that adapts to the source morphology, we sample the lens macro model once more, keeping the hyper-parameters of the source fixed at the highest-likelihood result.
@@
The Bayesian evidence naturally incorporates a penalty for increased model complexity – penalising models with additional free parameters, provides an objective basis for the principle of Occam’s razor. Typically, when comparing models using the difference in their log-evidences, $\Delta\ln\mathcal{Z}$, a difference in $\Delta\ln{\mathcal{Z}}>1$ is considered significant, $>2.5$ strong, and $>5$ is decisive, in favour of the model with the highest evidence.
@@ 4 Results
To disentangle the baryonic matter from the DM contribution and to reveal the radial arc, a model for the lens light distribution is essential. We fitted the lens light in both HST/WFC3 filters, F814W and F475W, using the MGE method described in Section 3.3. During the fitting process, the radial and tangential lensed images were masked to minimise contamination from the source planes. The PSF effects were incorporated by modelling the PSFr model as a sum of circular Gaussians. The F814W lens light model was used to trace the stellar distribution, as it is a better tracer of the stellar mass budged and has less contamination from the radial arc. For the lens modelling, we utilised the F475W image after the lens light subtraction. The left panel of Fig. 3 shows the F475W lens-subtracted image, and Table 1 presents the MGE decomposition for the F814W band.
Hereafter, all parameter estimates are derived from the final sampling chain. Unless otherwise stated, values represent the median of each parameter’s one-dimensional marginalised posterior probability distribution, with uncertainties corresponding to the $16^{\text{th}}$ and $84^{\text{th}}$ percentiles.
Table 1: MGE components of the HST/F814W image. The columns, from left to right, show the surface brightness of each Gaussian component, the MGE width, and the observed axis ratio. The MGE units were converted from counts to physical values following Trick et al. (2016), accounting for the redshift dimming effect and assuming a solar ST magnitude of $M_{\sun,\text{F814W}}=5.35$ from Willmer (2018).
@@ 4.1 Fiducial model
The fiducial model (M1) is composed by a stellar mass component, an elliptical NFW [^11] This is obtained by taking $\gamma_{\text{DM}}=1$ in Eq. 15. halo, and a central BH.
@@
The median of the one-dimensional marginalised posterior of the fiducial model and their uncertainties are summarised in Table 2. Fig. 4 presents the two-dimensional posterior distributions for the BH mass and other parameters that exhibit significant degeneracies with it.
@@
$\epsilon^{\text{sh}}_{1}$ $0.01^{+0.01}_{-0.01}$
$\epsilon^{\text{sh}}_{2}$ $0.05^{+0.01}_{-0.01}$
@@
Our fiducial model favours a BH mass of $\log_{10}(M_{\text{BH}}/M_{\sun})=10.56^{+0.07}_{-0.08}$, with a notable degeneracy with the mass-to-light ratio and the DM characteristic density, as shown in Fig. 4. However, no significant degeneracy was observed between the BH mass and the anisotropy parameter. Similarly, the Einstein mass within the Cosmic Horseshoe, $M_{\text{Ein}}$, does not exhibit a strong degeneracy with the BH mass.
@@
Figure 4: Two-dimensional posterior distributions for the parameters of the fiducial model. Only parameters that show a strong degeneracy with the BH mass are displayed. The inset plot on the top-right present the covariance between the BH mass and Einstein mass within the Cosmic Horseshoe ring. Contours are the $1$ and $2\sigma$ credible intervals, respectively.
We recovered a total projected mass within $R_{\text{Ein}}$ of $M_{\text{Ein}}=5.45^{+0.02}_{-0.03}\times 10^{12}M_{\sun}$, consistent with our prior and the EPL model, albeit slightly higher than the value reported by Schuldt et al. (2019). The DM fraction within $R_{e}$ was found to be $f_{\text{DM}}(\leq R_{e})=0.72^{+0.02}_{-0.02}$, which is also higher than Schuldt et al. (2019) and Spiniello et al. (2011), though it remains consistent with the $1\sigma$ range of the latter.
Fig. 5 shows the highest-likelihood lens model, alongside the normalised residuals, and the source plane reconstruction of source s1 at $z=1.961$. Our model successfully reproduces the radial arc and its counter-image, while reconstructing the source’s surface brightness, which exhibits an irregular morphology.
Figure 5: Highest-likelihood lens model under the fiducial configuration. Panels, from top left to bottom right, show the observed image, the lensed source model, the normalised residuals, and the source reconstruction. All images are in units of electrons per second. There is also an interpolated version of the source, if preferred.
@@ 4.2 Alternative mass models
To assess the robustness of our BH detection, we now investigate a range of alternative mass models. We first consider perturbations to our fiducial model and subsequently explore more flexible configurations. For each alternative model, we ran the full non-linear sampling again, following the steps outlined in section 3.5. Table 3 summarises the various mass models considered and their inferred BH masses, along with their corresponding Bayesian evidence relative to the fiducial model.
@@
Table 3: Joint fiducial and alternative mass models. The columns, from left to right, are: model name identification, mass model configuration, median of the BH mass posterior distribution, and the difference in the natural logarithm of the Bayesian evidences relative to the fiducial model (M1). Superscripts on the left side means the number of free components, i.e., ${}^{3}\beta_{\text{star}}$ implies three free anisotropy parameters.
Model ID Mass model $\log_{10}(M_{\text{BH}}/M_{\sun})$ $\Delta\ln{\mathcal{Z}}$
M1 $\Upsilon_{\star}$ + ell. NFW ($r_{s}=10R_{e}$) + $\beta_{\text{star}}$ + BH $10.56^{+0.07}_{-0.08}$ 0.00
M2 $\Upsilon_{\star}$ + ell. gNFW ($r_{s}=10R_{e}$) + $\beta_{\text{star}}$ + BH $10.57^{+0.07}_{-0.09}$0.53
M3 $\Upsilon_{\star}$ + ell. NFW ($r_{s}=10R_{e}$) + ${}^{3}\beta_{\text{star}}$ + BH $10.45^{+0.11}_{-0.14}$3.48
M4 ${}^{3}\Upsilon_{\star}$ + ell. NFW ($r_{s}=10R_{e}$) + $\beta_{\text{star}}$ + BH $10.53^{+0.10}_{-0.11}$ 2.01
M5 $\Upsilon_{\star}$ + ell. NFW + $\beta_{\text{star}}$ + BH $10.56^{+0.08}_{-0.08}$0.38
M6 $\Upsilon_{\star}$ + ell. NFW ($r_{s}=10R_{e}$) + $\beta_{\text{star}}$ + BH w/ cyl. velocity ellipsoids $10.55^{+0.08}_{-0.09}$0.36
M7 $\Upsilon_{\star}$ + ell. NFW ($r_{s}=10R_{e}$) + $\beta_{\text{star}}$ + BH w/ Delaunay pixelisation $10.55^{+0.08}_{-0.08}$2.27
M8 $\Upsilon_{\star}$ + ell. NFW ($r_{s}=10R_{e}$) + $\beta_{\text{star}}$ + BH w/o Horseshoe $10.51^{+0.07}_{-0.09}$ 0.38
M9 ${}^{3}\Upsilon_{\star}$ + ell. gNFW ($r_{s}=10R_{e}$) + ${}^{3}\beta_{\text{star}}$ + BH $10.50^{+0.10}_{-0.32}$ 8.29
M10 Gaussian $\Upsilon_{\star}$ + sph. NFW ($r_{s}=10R_{e}$) + ${}^{8}\beta_{\text{star}}$ + BH $10.55^{+0.10}_{-0.07}$ 1.17
M11 ${}^{3}\Upsilon_{\star}$ + ell. gNFW w/ main $c(M,z)$ + ${}^{3}\beta_{\text{star}}$ + BH $10.15^{+0.17}_{-0.30}$ 2.92
M12 ${}^{3}\Upsilon_{\star}$ + ell. gNFW w/ $1\sigma$ bellow $c(M,z)$ + ${}^{3}\beta_{\text{star}}$ + BH $10.33^{+0.07}_{-0.13}$ 10.73
M13 ${}^{3}\Upsilon_{\star}$ + ell. gNFW w/ $1\sigma$ above $c(M,z)$ + ${}^{3}\beta_{\text{star}}$ + BH $10.59^{+0.04}_{-0.10}$ 11.48
M14 $\Upsilon_{\star}$ + ell. NFW ($r_{s}=10R_{e}$) + $\beta_{\text{star}}$ 16.31
M15 ${}^{3}\Upsilon_{\star}$ + ell. NFW ($r_{s}=10R_{e}$) + $\beta_{\text{star}}$ 11.17
@@ 4.2.1 Fiducial model perturbations
In our fiducial model M1, we considered an NFW halo for which the inner density slope is fixed. However, a steeper DM density profile, in the inner regions, could potentially compensate for the BH mass, leading to a more cuspy rather than cored DM distribution. To explore this possibility, we considered an alternative model (M2) in which the DM halo is parameterised by a gNFW profile, allowing the inner density slope, $\gamma_{\text{DM}}$, to vary.
Fig. 6 shows the two-dimensional posterior distributions for the BH mass, the DM density slope, and the Einstein mass, with the median values of the fiducial model indicated by brown dashed lines. We can see that $\gamma_{\text{DM}}$ exhibits only a modest degeneracy with the BH mass and a marginal degeneracy with the Einstein mass. Furthermore, the DM inner density slope was found to be $\gamma_{\text{DM}}=1.06^{+0.05}_{-0.07}$, in agreement with an NFW profile. The BH mass is consistent with the fiducial model within $2\sigma$.
Figure 6: Two-dimensional posterior distributions for the parameters of model M2, which differs from the fiducial model by assuming a gNFW halo. Only the inner DM density slope, the BH mass, and the Einstein mass are displayed. The brown dashed lines shows the posterior median of the fiducial model for comparison. Contours are the $1$ and $2\sigma$ credible intervals, respectively.
@@
To assess the impact of stellar anisotropy on the BH mass estimate, we introduced a stellar anisotropy profile, $\beta_{\text{star}}(r)$, in model M3. This profile is constructed using the luminous MGE components (Cappellari, 2008) to define regions with distinct anisotropy parameters. We used the Gaussian width as a proxy for the radius of influence of the parameter. Specifically:
@@
In Fig. 7, we present the stellar anisotropy profile for model M3, calculated following Cappellari (2008). Unlike other BCGs hosting central SMBHs, the anisotropy profile of our system near the galaxy centre is qualitatively distinct, showing no dominance of tangential orbits. (e.g., Thomas et al., 2014; Mehrgan et al., 2019). It is worth noting, however, that we cannot resolve the galaxy core (if present), where tangential orbits are typically expected. Moreover, the uncertainties near the centre suggest that $\beta_{\text{star}}$ could also assume negative values in this region. Despite that, the radial variation observed in our profile is broadly consistent with findings by Gerhard et al. (2001) in elliptical galaxies and with simulations of massive ETGs (Wu et al., 2014). The BH mass derived from this model is slightly lower than the value obtained from the fiducial model but remains within the $1\sigma$ confidence level.
Figure 7: M3 anisotropy profile. The solid line shows the orbital anisotropy profile of the stars, and the gray band represents the $1\sigma$ credible band. The brown point is the constant anisotropy inferred by the fiducial model M1, with its associated error bar. The horizontal dashed line corresponds to the isotropic case.
@@
Another way to account for the BH mass is by increasing the stellar mass at the galaxy’s centre, representing an excess mass linked to a gradient in the stellar mass-to-light ratio (e.g., Smith et al., 2017a). To test this hypothesis, in model M4, we allowed the mass-to-light ratio to vary across different sets of Gaussian components. As in model M3, we defined the mass-to-light ratio for each luminous Gaussian component based on the Gaussian width, applying the same width constraints. This setup resulted in three distinct mass-to-light ratios ($\Upsilon^{0}_{\star},\Upsilon^{1}_{\star},\Upsilon^{2}_{\star}$), which were constrained during the non-linear sampling to follow a decreasing gradient with radius.
In Fig. 8, we present the two-dimensional posterior distributions for the BH mass and the three mass-to-light ratios in model M4. The top-right inset plot compares the mass-to-light ratio profiles of M1 and M4. The gradient profile of model M4 is shown in black, while the constant profile of the fiducial model is shown in brown. Although model M4 suggests a radial gradient, the large uncertainties make it consistent with the constant value recovered by the fiducial model. By this figure, it is clear the strong degeneracy between these parameters. Despite that, the recovered BH mass is consistent with the fiducial model within $1\sigma$.
Figure 8: Two-dimensional posterior distributions for the parameters of model M4, which differs from the fiducial model by a gradient mass-to-light ratio. Only the three mass-to-light ratios, the BH mass, and the Einstein mass are displayed. The brown dashed lines shows the posterior median of the fiducial model for comparison. Contours are the $1$ and $2\sigma$ credible intervals, respectively. The top-right inset shows the projected radial stellar mass-to-light profiles for the fiducial model (brown) and model M4 in black. The shaded regions are the $1\sigma$ level.
@@
In the fiducial model, we assume a DM scale radius fixed at ten times the galaxy’s effective radius, and although motivated by simulations, such rigid constraints could bias the results. To evaluate the impact of this assumption, in model M5 we allowed the scale radius to vary. This model recovered a scale radius of $r_{s}=19.16^{+2.30}_{-0.97}$, interestingly consistent with the fiducial model’s assumption. Moreover, the BH mass remains consistent with the fiducial value within $1\sigma.$

### v3

@@ 3.2 Dynamical modelling
and applying the boundary condition $\nu\overline{v^{2}_{r}}=0$ as $r\rightarrow 0$, the Jeans Eqs. 7 simplify to:
@@
For this purpose, we employed the Jeans Anisotropic Modelling (JAM) method (Cappellari, 2008, 2020), as implemented in the Jampy software, to perform the stellar dynamical modelling. This approach assumes that the galaxy’s mass distribution can be parameterised as a sum of concentric elliptical Gaussians (see Section 3.3).
@@ 3.3 Multi-Gaussian Expansion (MGE)
The stellar light and mass component are modelled using the MGE method (Emsellem et al., 1994; Cappellari, 2002), which parametrises the stellar surface brightness as a sum of concentrict elliptical Gaussians with the same orientation. If $I(x^{\prime},y^{\prime})$ represents the stellar surface brightness, its MGE parametrisation is given by
@@
Here, $M_{j}=\Upsilon_{\star}L_{j}$ represents the mass of the $j^{\text{th}}$ Gaussian component, with $L_{j}$ luminosity and $\sigma_{j}$ dispersion. The deprojected axial ratio, $q_{j}$, is given by
@@ 3.4 Mass profile
We described the mass profile of the main deflector using the MGE method, assuming a multicomponent mass model composed by (i) a stellar mass component; (ii) a DM halo component; and (iii) an additional central mass concentration representing an SMBH.
@@
where $\rho_{s}$ is a characteristic density at the scale radius $r_{s}$, and $\gamma_{\text{DM}}$ is the inner density slope that allows the profile to be cuspier ($\gamma_{\text{DM}}>1$) or cored ($\gamma_{\text{DM}}=0$). For $\gamma_{\text{DM}}=1$, the profile reduces to the classical NFW (Navarro et al., 1997). To include this halo contribution in the dynamical model, we followed Cappellari et al. (2013) and also parametrised the DM component using the MGE method [^11] This parametrisation is performed iteratively, starting with 20 Gaussian components and adding more if the absolute deviation from the target analytical profile exceeds 10% in any radial bin. The process continues until all deviations fall below the threshold or a maximum of 45 components is reached. Based on our tests, 20 Gaussians are typically sufficient to accurately represent NFW or gNFW profiles.. Finally, the SMBH is modelled as an additional Gaussian with a small scale radius ($\sigma=0.01\arcsec$).
@@ 3.5 Joint modelling
We break the modelling process into the following stages, so as to make it more tractable
@@
    Dynamical Model Fitting: We begin by fitting only the dynamical model, but with the addition of the prior term for  $M_{\text{Ein}}$ and the conjugate regions of s1 image. This provides an initial estimate of the mass parameters.
@@
    Mass Parameter Resampling: After obtaining the source grid, we resample the mass parameters using the full (lensing+dynamics) likelihood, with s1 reconstructed based on the highest-likelihood model from the previous chain. This step improves the estimate of the mass parameters.
@@
    Source Plane Resampling: Using the highest-likelihood model of the previous stage, we resample the parameters describing the source plane pixelisation and regularisation. The source is now reconstructed in the source plane using a brightness-adaptive Voronoi mesh grid, with Natural Neighbour interpolation (Sibson, 1981) and a regularisation scheme that adapts the degree of smoothing based on the source’s surface brightness. This ensures that brighter regions are reconstructed at higher resolution, while regions with lower SNR are more regularised. In PyAutoLens notation, this corresponds to a KMeans mesh grid, with VoronoiNN pixelisation, and AdaptiveBrightnessSplit regularisation.
@@
    Final Resampling of the Lens Macro Model: Finally, using the source grid that adapts to the source morphology, we sample the lens macro model once more, keeping the hyper-parameters of the source fixed at the highest-likelihood result of the previous step.
@@
The Bayesian evidence naturally incorporates a penalty for increased model complexity – penalising models with additional free parameters-, and provides an objective basis for the principle of Occam’s razor. Typically, when comparing models using the difference in their log-evidences, $\Delta\ln\mathcal{Z}$, a difference in $\Delta\ln{\mathcal{Z}}>1$ is considered significant, $>2.5$ strong, and $>5$ is decisive, in favour of the model with the highest evidence (e.g. Jeffreys, 1961).
@@ 4 Results
To disentangle the baryonic matter from the DM contribution and to reveal the radial arc, a model for the lens light distribution is essential. We fitted the lens light in both HST/WFC3 filters, F814W and F475W, using the MGE method described in Section 3.3. During the fitting process, the radial and tangential lensed images were masked to minimise contamination from the source planes. The PSF effects were incorporated by modelling the PSFr model as a sum of circular Gaussians. The F814W lens light model was used to trace the stellar distribution, as it is a better tracer of the stellar mass budged and has less contamination from the radial arc. For the lens modelling, we utilised the F475W image after the lens light subtraction, as the radial source is better seem in this band. The left panel of Fig. 3 shows the F475W lens-subtracted image, and Table 1 presents the MGE decomposition for the F814W band.
In the following sections, we discuss our fiducial model and explore various mass models to assess potential systematic effects on the SMBH mass inference. Table 3 summarises these models and their resulting $M_{\text{BH}}$ values. Unless otherwise, all parameter estimates are derived from the final sampling chain, with reported values representing the median of each parameter’s one-dimensional marginalised posterior distribution, with uncertainties corresponding to the $16^{\text{th}}$ and $84^{\text{th}}$ percentiles.
Table 1: MGE components of the HST/F814W image. The columns, from left to right, show the surface brightness of each Gaussian component, the MGE width, and the observed axis ratio. The MGE units were converted from counts to physical values following Trick et al. (2016), accounting for the redshift dimming effect and assuming a solar Space Telescope magnitude (STmag) of $M_{\sun,\text{F814W}}=5.35$ from Willmer (2018).
@@ 4.1 Fiducial model
The fiducial model (M1) is composed by a stellar mass component, an elliptical NFW [^12] This is obtained by taking $\gamma_{\text{DM}}=1$ in Eq. 15. halo, and a central SMBH.
@@
The median of the one-dimensional marginalised posterior of the fiducial model and their uncertainties are summarised in Table 2. Fig. 4 presents the two-dimensional posterior distributions for the SMBH mass and other parameters that exhibit significant degeneracies with it.
@@
$\epsilon^{\text{sh}}_{1}$ $-0.01^{+0.01}_{-0.01}$
$\epsilon^{\text{sh}}_{2}$ $-0.05^{+0.01}_{-0.01}$
@@
Our fiducial model favours an SMBH mass of $\log_{10}(M_{\text{BH}}/M_{\sun})=10.56^{+0.07}_{-0.08}$, therefore an UMBH. The SMBH mass shows a notable degeneracy with the mass-to-light ratio and the DM characteristic density, as shown in Fig. 4. However, no significant degeneracy was observed between the SMBH mass and the anisotropy parameter. Similarly, the Einstein mass within the Cosmic Horseshoe, $M_{\text{Ein}}$, does not exhibit a strong degeneracy with $M_{\text{BH}}$ top-right inset plot in Fig. 4.
@@
Figure 4: Two-dimensional posterior distributions for the parameters of the fiducial model. Only parameters that show a strong degeneracy with the SMBH mass are displayed. The inset plot on the top-right present the covariance between the SMBH mass and Einstein mass within the Cosmic Horseshoe ring. Contours are the $1$ and $2\sigma$ credible intervals, respectively.
We recovered a total projected mass within $R_{\text{Ein}}$ of $M_{\text{Ein}}=5.45^{+0.02}_{-0.03}\times 10^{12}M_{\sun}$, consistent with our prior and the EPL model, albeit slightly higher than the value reported by Dye et al. (2008) and Schuldt et al. (2019). The DM fraction within $R_{e}$ was found to be $f_{\text{DM}}(\leq R_{e})=0.72^{+0.02}_{-0.02}$, which is also higher than Schuldt et al. (2019) and Spiniello et al. (2011), though it remains consistent within the $1\sigma$ range.
Fig. 5 shows the highest-likelihood lens model, alongside the normalised residuals and the source plane reconstruction of source s1 at $z_{s1}=1.961$. Our model successfully reproduces the radial arc and its counter-image, while reconstructing the source’s surface brightness, which exhibits an irregular morphology.
Figure 5: Highest-likelihood lens model under the fiducial configuration. Panels, from top left to bottom right, show the observed image, lensed source model, normalised residuals, and source reconstruction. All images are in units of electrons per second.
@@ 4.2 Alternative mass models
To assess the robustness of our SMBH detection, we now investigate a range of alternative mass models. We first consider perturbations to our fiducial model and subsequently explore more flexible configurations. For each alternative model, we ran the full non-linear sampling again, following the steps outlined in Section 3.5. Table 3 summarises the various mass models considered and their inferred SMBH masses, along with their corresponding Bayesian evidence relative to the fiducial model.
@@
Table 3: Joint fiducial and alternative mass models. The columns, from left to right, are: model name identification, mass model configuration, median of the SMBH mass posterior distribution, the difference in the natural logarithm of the Bayesian evidences relative to the fiducial model (M1), section where the model is discussed, and the total number of free parameters in the mass model. Superscripts on the left side means the number of free components, i.e., ${}^{3}\beta_{\text{star}}$ implies three free anisotropy parameters.
Model ID Mass model $\log_{10}(M_{\text{BH}}/M_{\sun})$ $\Delta\ln{\mathcal{Z}}$ Section $N_{\text{par}}$
M1 $\Upsilon_{\star}$ + ell. NFW ($r_{s}=10R_{e}$) + $\beta_{\text{star}}$ + BH $10.56^{+0.07}_{-0.08}$ 0.00 4.1 8
M2 $\Upsilon_{\star}$ + ell. gNFW ($r_{s}=10R_{e}$) + $\beta_{\text{star}}$ + BH $10.57^{+0.07}_{-0.09}$0.53 4.2.1 9
M3 $\Upsilon_{\star}$ + ell. NFW ($r_{s}=10R_{e}$) + ${}^{3}\beta_{\text{star}}$ + BH $10.45^{+0.11}_{-0.14}$3.48 4.2.1 10
M4 ${}^{3}\Upsilon_{\star}$ + ell. NFW ($r_{s}=10R_{e}$) + $\beta_{\text{star}}$ + BH $10.53^{+0.10}_{-0.11}$ 2.01 4.2.1 10
M5 $\Upsilon_{\star}$ + ell. NFW + $\beta_{\text{star}}$ + BH $10.56^{+0.08}_{-0.08}$0.38 4.2.1 9
M6 $\Upsilon_{\star}$ + ell. NFW ($r_{s}=10R_{e}$) + $\beta_{\text{star}}$ + BH w/ cyl. velocity ellipsoids $10.55^{+0.08}_{-0.09}$0.36 4.2.1 8
M7 $\Upsilon_{\star}$ + ell. NFW ($r_{s}=10R_{e}$) + $\beta_{\text{star}}$ + BH w/ Delaunay pixelisation $10.55^{+0.08}_{-0.08}$2.27 4.2.1 8
M8 $\Upsilon_{\star}$ + ell. NFW ($r_{s}=10R_{e}$) + $\beta_{\text{star}}$ + BH w/o Horseshoe $10.51^{+0.07}_{-0.09}$ 0.38 4.2.1 8
M9 ${}^{3}\Upsilon_{\star}$ + ell. gNFW ($r_{s}=10R_{e}$) + ${}^{3}\beta_{\text{star}}$ + BH $10.50^{+0.10}_{-0.32}$ 8.29 4.2.2 13
M10 Gaussian $\Upsilon_{\star}$ + sph. NFW ($r_{s}=10R_{e}$) + ${}^{8}\beta_{\text{star}}$ + BH $10.55^{+0.10}_{-0.07}$ 1.17 4.2.2 16
M11 ${}^{3}\Upsilon_{\star}$ + ell. gNFW w/ main $c(M,z)$ + ${}^{3}\beta_{\text{star}}$ + BH $10.15^{+0.17}_{-0.30}$ 2.92 4.2.2 13
M12 ${}^{3}\Upsilon_{\star}$ + ell. gNFW w/ $1\sigma$ below $c(M,z)$ + ${}^{3}\beta_{\text{star}}$ + BH $10.33^{+0.07}_{-0.13}$ 10.73 4.2.2 13
M13 ${}^{3}\Upsilon_{\star}$ + ell. gNFW w/ $1\sigma$ above $c(M,z)$ + ${}^{3}\beta_{\text{star}}$ + BH $10.59^{+0.04}_{-0.10}$ 11.48 4.2.2 13
M14 $\Upsilon_{\star}$ + ell. NFW ($r_{s}=10R_{e}$) + $\beta_{\text{star}}$ 16.31 4.3 7
M15 ${}^{3}\Upsilon_{\star}$ + ell. NFW ($r_{s}=10R_{e}$) + $\beta_{\text{star}}$ 11.17 4.3 9
@@ 4.2.1 Fiducial model perturbations
In our fiducial model M1, we considered an NFW halo for which the inner density slope is fixed. However, a steeper DM density profile, in the inner regions, could potentially compensate for the SMBH mass, leading to a more cuspy rather than cored DM distribution. To explore this possibility, we considered an alternative model (M2) in which the DM halo is parameterised by a gNFW profile, allowing the inner density slope, $\gamma_{\text{DM}}$, to vary.
Fig. 6 shows the two-dimensional posterior distributions for the SMBH mass, the DM density slope, and the Einstein mass, with the median values of the fiducial model indicated by brown dashed lines. We can see that $\gamma_{\text{DM}}$ exhibits only a modest degeneracy with $M_{\text{BH}}$ and a marginal degeneracy with the Einstein mass. Furthermore, the DM inner density slope was found to be $\gamma_{\text{DM}}=1.06^{+0.05}_{-0.07}$, in agreement with an NFW profile. The SMBH mass is consistent with the fiducial model within $2\sigma$.
Figure 6: Two-dimensional posterior distributions for the parameters of model M2, which differs from the fiducial model by assuming a gNFW halo. Only the inner DM density slope, the SMBH mass, and the Einstein mass are displayed. The brown dashed lines shows the posterior median of the fiducial model for comparison. Contours are the $1$ and $2\sigma$ credible intervals, respectively.
@@
To assess the impact of stellar anisotropy on $M_{\text{BH}}$, we introduced a stellar anisotropy profile, $\beta_{\text{star}}(r)$, in model M3. This profile is constructed using the luminous MGE components (Cappellari, 2008) to define regions with distinct anisotropy parameters. We used the Gaussian width as a proxy for the radius of influence of the parameter. Specifically:
@@
In Fig. 7, we present the stellar anisotropy profile for model M3, calculated following Cappellari (2008). Unlike other BCGs hosting central SMBHs, the anisotropy profile of our system near the galaxy centre is qualitatively distinct, showing no dominance of tangential orbits (e.g., Thomas et al., 2014; Mehrgan et al., 2019). It is worth noting, however, that we cannot resolve the galaxy core (if present), where tangential orbits are typically expected. Moreover, the uncertainties near the centre suggest that $\beta_{\text{star}}$ could also assume negative values in this region. Despite that, the radial variation observed in our profile is broadly consistent with findings by Gerhard et al. (2001) in elliptical galaxies and with simulations of massive ETGs (Wu et al., 2014). The SMBH mass derived from this model is slightly lower (see Table 3) than the value obtained from the fiducial model, but remains within the $1\sigma$ confidence level.
Figure 7: M3 anisotropy profile. The solid line shows the orbital anisotropy profile of the stars, and the grey band represents the $1\sigma$ credible band. The brown point is the constant anisotropy inferred by the fiducial model M1, with its associated error bar. The horizontal dashed line corresponds to the isotropic case.
@@
Another way to account for the SMBH mass is by increasing the stellar mass at the galaxy’s centre, representing an excess mass linked to a gradient in the stellar mass-to-light ratio (e.g., Smith et al., 2017a). To test this hypothesis, in model M4, we allowed the mass-to-light ratio to vary across different sets of Gaussian components. As in model M3, we defined the mass-to-light ratio for each luminous Gaussian component based on the Gaussian width, applying the same width constraints. This setup resulted in three distinct mass-to-light ratios ($\Upsilon^{0}_{\star},\Upsilon^{1}_{\star},\Upsilon^{2}_{\star}$), which were constrained during the non-linear sampling to follow a decreasing gradient with radius.
In Fig. 8, we present the two-dimensional posterior distributions for $M_{\text{BH}}$ and the three mass-to-light ratios in model M4. The top-right inset plot compares the mass-to-light ratio profiles of M1 and M4. The gradient profile of model M4 is shown in black, while the constant profile of the fiducial model is shown in brown. Although model M4 suggests a radial gradient, the large uncertainties make it consistent with the constant value recovered by the fiducial model. By this figure, it is clear the strong degeneracy between these parameters. Despite that, the recovered value of $\log_{10}(M_{\text{BH}}/M_{\sun})=10.53^{+0.10}_{-0.11}$, is consistent with the fiducial model within $1\sigma$.
Figure 8: Two-dimensional posterior distributions for the parameters of model M4, which differs from the fiducial model by a gradient mass-to-light ratio. Only the three mass-to-light ratios, the SMBH mass, and the Einstein mass are displayed. The brown dashed lines shows the posterior median of the fiducial model for comparison. Contours are the $1$ and $2\sigma$ credible intervals, respectively. The top-right inset shows the projected radial stellar mass-to-light profiles for the fiducial model (brown) and model M4 in black. The shaded regions are the $1\sigma$ level.
@@
In the fiducial model, we assume a DM scale radius fixed at ten times the galaxy’s effective radius, and although motivated by simulations, such rigid constraints could bias the results. To evaluate the impact of this assumption, in model M5 we allowed the scale radius to vary. This model recovered a scale radius of $r_{s}=19.16^{+2.30}_{-0.97}$ arcsec, interestingly consistent with the fiducial model’s assumption ($r_{s}=21$ arcsec). Moreover, the SMBH mass (see Table 3) remains consistent with the fiducial value within $1\sigma.$